OpenStudy (anonymous):

solve the equation log(x-5) + log(x+2) = 1 algebraically . find any extraneous solution. if there is no extaneouse answer why?

OpenStudy (kira_yamato):

log(x-5) + log(x+2) = 1 log[(x-5)(x+2)] = 1 (x-5)(x+2) = 10 x^2 - 3x - 20 = 0 x = 6.216990566028302 or -3.2169905660283016 (N.A. since x > 0) x ≈ 6.217